40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat May 16 21:05:24 2015, rana, Update, General, some status 48.png39.png
    Reply  Mon May 18 09:50:00 2015, ericq, Update, General, some status 
       Reply  Mon May 18 11:59:07 2015, rana, Update, General, some status 
          Reply  Tue May 19 16:18:57 2015, ericq, Update, General, crons fixed 
    Reply  Mon May 18 16:28:18 2015, ericq, Update, General, some status 
       Reply  Mon May 18 18:03:12 2015, rana, Update, General, some status CPUtrend.png
          Reply  Tue May 19 00:19:23 2015, rana, Update, General, some status Screen_Shot_2015-05-19_at_12.17.39_AM.png
             Reply  Wed May 20 03:08:27 2015, rana, Update, General, some status 
    Reply  Mon May 18 17:42:14 2015, rana, ericQ, Update, General, some status 
       Reply  Tue May 19 18:55:12 2015, rana, ericQ, Update, General, some status 
          Reply  Wed May 20 11:41:59 2015, ericq, Update, General, some status 
    Reply  Mon Sep 21 00:51:36 2015, rana, Update, General, op340m, autoburt cron =? megatron 
       Reply  Mon Sep 21 11:40:30 2015, ericq, Update, General, Megatron maitenence 
          Reply  Tue Dec 1 17:26:14 2015, Koji, Update, General, Megatron maitenence 
             Reply  Tue Dec 1 20:20:16 2015, Koji, Update, General, Megatron maitenence 
Message ID: 11294     Entry time: Sat May 16 21:05:24 2015     Reply to this: 11297   11301   11303   11624
Author: rana 
Type: Update 
Category: General 
Subject: some status 

1) Checked the N2 pressures: the unregulated cylinder pressures are both around 1500 PSI. How long until they get to 1000?

2) The IMC has been flaky for a day or so; don't know why. I moved the gains in the autolocker so now the input gain slider to the MC board is 10 dB higher and the output slider is 10 dB lower. This is updated in the mcdown and mcup scripts and both committed to SVN. The trend shows that the MC was wandering away after ~15 minutes of lock, so I suspected the WFS offsets. I ran the offsets script (after flipping the z servo signs and adding 'C1:' prefix). So far powers are good and stable.

3) pianosa was unresponsive and I couldn't ssh to it. I powered it off and then it came back.

4) Noticed that DAQD is restarting once per hour on the hour. Why?

5) Many (but not all) EPICS readbacks are whiting out every several minutes. I remote booted c1susaux since it was one of the victims, but it didn't change any behavior.

6) The ETMX and ITMX have very different bounce mode response: should add to our Vent Todo List. Double checked that the bounce/roll bandstop is on and at the right frequency for the bounce mode. Increased the stopband from 40 to 50 dB to see if that helps.

7) op340 is still running ! The only reason to keep it alive is its crontab:

op340m:SUS>crontab -l

07 * * * * /opt/rtcds/caltech/c1/burt/autoburt/burt.cron >> /opt/rtcds/caltech/c1/burt/burtcron.log
#46 * * * * /opt/rtcds/caltech/c1/scripts/general/scripto_cron /opt/rtcds/caltech/c1/scripts/PSL/FSS/FSSSlowServo > /cvs/cds/caltech/logs/scripts/FSSslow.cronlog 2>&1
#14,44 * * * * /cvs/cds/caltech/conlog/bin/check_conlogger_and_restart_if_dead
15,45 * * * * /opt/rtcds/caltech/c1/scripts/SUS/rampdown.pl > /dev/null 2>&1
#10 * * * *  /opt/rtcds/caltech/c1/scripts/general/scripto_cron /opt/rtcds/caltech/c1/scripts/MC/autolockMCmain40m >/cvs/cds/caltech/logs/scripts/mclock.cronlog 2>&1
#27 * * * * /opt/rtcds/caltech/c1/scripts/general/scripto_cron /opt/rtcds/caltech/c1/scripts/PSL/FSS/RCthermalPID.pl >/cvs/cds/caltech/logs/scripts/RCthermalPID.cronlog 2>&1

00 0 * * * /var/scripts/ntp.sh > /dev/null 2>&1
#00 4 * * * /opt/rtcds/caltech/c1/scripts/RGA/RGAlogger.cron >> /cvs/cds/caltech/users/rward/RGA/RGAcron.out 2>&1
#00 6 * * * /cvs/cds/scripts/backupScripts.pl
00 7 * * * /opt/rtcds/caltech/c1/scripts/AutoUpdate/update_conlog.cron
00 8 * * * /opt/rtcds/caltech/c1/scripts/crontab/backupCrontab

added a new script (scripts/SUS/rampdown.py) which decrements every 30 minutes if needed. Added this to the megatron crontab and commented out the op340m crontab line. IF this works for awhile we can retire our last Solaris machine.

8) To see if we could get rid of the wandering PCDRIVE noise, I looked into the NPRO temperatures: was - T_crystal = 30.89 C, T_diode1 = 21 C, T_diode2 = 22 C. I moved up the crystal temp to 33.0 C, to see if it could make the noise more stable. Then I used the trimpots on the front of the controller to maximimize the laseroutput at these temperatures; it was basically maximized already. Lets see if there's any qualitative difference after a week. I'm attaching the pinout for the DSUB25 diagnostics connector on the back of the box. Aidan is going to help us record this stuff with AcroMag tech so that we can see if there's any correlation with PCDRIVE. The shifts in FSS_SLOW coincident with PCDRIVE noise corresponds to ~100 MHz, so it seems like it could be NPRO related.

 

Attachment 1: 48.png  82 kB  Uploaded Sat May 16 23:52:43 2015  | Hide | Hide all
48.png
Attachment 2: 39.png  525 kB  Uploaded Sun May 17 01:31:19 2015  | Hide | Hide all
39.png
ELOG V3.1.3-